# arXiv:2603.00269v1[stat.ME] 27 Feb 2026

### v1

@@ 4.1 Simulation Settings
To investigate $t$ regression on data with outliers, we considered a linear model with true $\beta=(1,5,10)$ and $n=100$. Additionally, we also considered data with outliers under high dimensional setting, generated using a linear model with $p=80$, true $\beta=\boldsymbol{0}_{p}$, and $n=300$. The covariates and errors were generated from the standard normal distribution, with contamination (i) $N(0,9)$, (ii) $\chi^{2}(4)-4$, (iii) $t(2)$, or (iv) two-point contamination in which any response value may become 5-5 or 5 with a probability $\lambda/2$ each. The contamination rates considered are $\lambda=$ 10%, 20%, and 30%.
@@ 4.2 Results when $p=3$
When the data is contaminated with $N(0,9)$ errors (see Figure 6) , although the $\nu$-estimation approaches do not necessarily give the lowest overall RMSE, they are able to give relatively low RMSE compared to OLS and Huber regression. As the contamination rate increases, the $\nu$-estimation approaches outperform most of the fixed $\nu$ approach estimations as they are successful in estimating $\nu$ close to truth. Similar pattern is shown in data with two-point contamination.
@@
[…]
@@ 4.3 Results when $p=80$
In high dimension setting, when the data is contaminated with $N(0,9)$ (see Figure 8), $\chi^{2}(4)-4$, or $t(2)$ errors, although $\hat{\nu}_{adj}$ approach does not necessarily give the lowest overall RMSE, it is able to give relatively lower RMSE compared to OLS and Huber regression. This may be explained by the fact that $\hat{\nu}_{adj}$ are closer the the truth compared to other $\nu$ estimation methods (and fixing large $\nu$) in high dimension setting.
@@
For data contaminated with two-point errors (see Figure 9), $\hat{\nu}_{adj}$ approach outperforms Huber and OLS when the contamination rate is low (e.g. 10%). As the contamination rate increases, Huber regression performs better as $\hat{\nu}_{adj}$ are overestimated, and hence the corresponding $\beta$ estimation performance is similar to fixing $\nu=30$. Nonetheless, the $\nu$-estimation approaches outperform OLS when the contamination rate is 20%. As the contamination rate increases to 30%, all $\nu$-estimation approaches perform as bad as most fixed-$\nu$ approaches and OLS. This may be caused by high contamination of outliers with a ”uniform” pattern in the data that t distribution fails to capture.
These results collectively suggest that the $\nu$-estimation approaches, and $\hat{\nu}_{adj}$ in some high dimension settings, offer a robust performance across various contamination types and levels, often outperforming classical regression methods.
@@ 5 Conclusions and Discussion
In this paper, we studied robust regression with Student’s t error and addressed the commonly asked question on how to choose the degree of freedom parameter in Student’s t distribution in practice. We compared 4 frequentist and Bayesian methods in estimating the degree of freedom $\nu$, and studied how these choices propagate to estimation of the regression coefficients $\beta$ in uncontaminated and contaminated data. In terms of $\beta$ estimation performance, we suggest estimating $\nu$ with adjusted profile likelihood when the $p/n$ ratio is not too large as it is the most robust in both heavy-tailed and light-tailed distribution, as well as in contaminated data and large $p$ (given moderate $n$) settings. However, to ensure stable $\nu$ and $\beta$ estimations, our proposed $\nu$-estimation methods require $p$ to be not too large relative to $n$. This is because high dimensionality may lead to larger residuals magnitude and tend to overestimate $\nu$. Therefore, when $p/n$ ratio is very large (it would be better if we were able to quantify this), optimizing the likelihood by fixing $\nu$ at some large values may be more efficient.
Appendix A Supplementary Material
Appendix B Derivation of observed Fisher information for $\beta$ and $\sigma$
@@
Appendix C Proof of Invariance Property
@@ C.1 Profile likelihood
$l^{P}(\nu;y^{\prime},x)$ $=C(\nu)-\frac{\nu+1}{2}\sum_{i=1}^{n}log\{\nu+z_{i}^{2}\}-n\log(a\hat{\sigma}_{\nu})=l^{P}(\nu;y,x)-n\log(a),$
@@
C.2 Adjusted Profile likelihood
@@
C.3 Full Independence Jeffreys
@@ C.4 Marginal Jeffrey’s
Appendix D More details on $\nu$-optimization
D.1 Out of bound values
@@
D.2 Choice of sample sizes and starting values
Fonseca et al. (2008) noted that, as a function of $\nu$, the profile log-likelihood function will increase without bound when $\sum_{i=1}^{n}(\hat{z_{i}}^{2}-1)^{2}<2n$ where $\hat{z_{i}}=(y_{i}-x_{i}^{T}\hat{\beta)}/\hat{\sigma}$, $\hat{\beta},\hat{\sigma}$ are the OLS estimators. Hence, to ensure that the optimization algorithm would converge, we first searched for large enough $n$ such that Fonseca et al. (2008)’s condition is violated. Next, we searched for $n$ such that all 5 objective functions do not have flatness issue by manually plotting out the functions for 20 simulated datasets. When we ran simulations using smaller $n$, for instance, $\nu=2$; $p=40$ using $n=100$, around 3.33% of the simulations had flatness issue and over 20% $\hat{\nu}_{adj}$ from the remaining successful converged simulations were greater than $10$, which are bad estimates. When $p$ was further increased to $60$, the flatness issue appeared in 55% of the simulations, and over 18% $\hat{\nu}_{adj}$ from the remaining successful converged simulations were greater than $10$.
@@ D.2 Choice of sample sizes and starting values
When there are no flatness issue (i.e., the simulation settings we consider), setting the initial value at true $\omega$ that was used to generate the data optimizes the optimization time since we expect the returned $\hat{\omega}$ to be close to the truth. In our simulation settings, using non-truth value as initial guess would still yield to stable estimates and RMSE results under the adjusted likelihood approach. The other 4 approaches tend to underestimate $\nu$, such phenomenon becomes more obvious as we set $\omega_{\text{init}}$ to be lower than the truth. For the purpose to compare the performances between the 5 approaches, the overall RMSE conclusion were consistent with results obtained from setting the initial guess at truth, in which we concluded that the adjusted likelihood approach resulted in lowest RMSE. For example, when we ran simulations on $n=300$; $\nu=2$; $p=60$ using the $1/5$ as initial guess of $\omega$, RMSE of $\hat{\nu}_{adj}$ is $0.5824363$, which is very close to that from setting initial guess at $1/2$ (i.e., truth). The other approaches yield RMSE at around 1.361.391.36-1.39, which are slightly inflated from setting initial guess at truth which gave 1.251.291.25-1.29 RMSE. If the initial guess of $\omega$ is set at $1/10$, RMSE of $\hat{\nu}_{adj}$ is $0.5860144$ and the other approaches yield RMSE at around 1.441.491.44-1.49. The boxplots below display the spread of $\hat{\nu}$ in each setting.
@@
Appendix E Extra Results of $\nu$ estimation simulations
@@
Appendix F Order of Bias and Variance for $\hat{\nu}$
@@
For the profile log-likelihood estimation of $\hat{\nu}$, both the variance and the bias has order $\mathcal{O}(n^{-1})$.
@@
From Reid (2013), asymptotically, the profile log-likelihood is
$j_{p}^{1/2}(\hat{\nu})\,(\hat{\nu}-\nu)\;\xrightarrow{\mathcal{L}}\;\mathcal{N}(0,I),$ (9)
@@
Define $f(\nu,\hat{\lambda}_{\nu})=\partial\ell(\nu,\hat{\lambda}_{\nu};y,x)/\partial\hat{\lambda}_{\nu}=0:=\ell_{\lambda}$. If $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\hat{\lambda}_{\nu}$ is invertiable, taking the derivative and set $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\nu=0$
$\frac{\partial\hat{\lambda}_{\nu}}{\partial\nu}=-\left(\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\hat{\lambda}_{\nu}}\right)^{-1}\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\hat{\lambda}_{\nu}}.$
@@
$j_{p}(\hat{\nu})=-\ell_{\nu\nu}+\ell_{\nu\lambda}(\ell_{\lambda\lambda})^{-1}\ell_{\lambda\nu}.$
@@
The term $\ell_{\lambda\lambda}$ is the observed Fisher information matrix for the adjusted profile likelihood, $j_{\lambda\lambda}$ (LABEL:eq:obsfisher), which has order of $\mathcal{O}(n)$. Therefore, $j_{p}(\hat{\nu})=\mathcal{O}(n)$,so we have the $Var(\hat{\nu})=\mathcal{O}(n^{-1})$.
Next, we analyze the bias part, we assume the maximum likelihood estimate of $\beta,\sigma$ and $\nu$ exists. The maximum profile likelihood estimates coincides with the maximum likelihood estimates, so it suffices to show the order of bias for the maximum likelihood estimate of $\nu$. We follow the proof by Cox and Snell (1968). The maximum likelihood equation $\ell^{\prime}(\hat{\nu},\hat{\beta},\hat{\sigma};y,x)=0$ to the first order for $\nu$ is
@@
$U^{i}=\frac{\partial\log\ell(\nu,\hat{\beta},\hat{\sigma};y,x)}{\partial\nu},V^{i}=\frac{\partial\log\ell(\nu,\hat{\beta},\hat{\sigma};y,x)}{\partial\nu^{2}},$
@@
To obtain a more refined answer, we replace it by the second-order equation
@@
$cov(\hat{\nu}-\nu,\ell^{\prime\prime}(\nu))=cov\left(\frac{\sum_{i=1}^{n}U^{i}}{I},V\right)=\frac{1}{I}cov\left(\sum_{i=1}^{n}U^{i},\sum_{i=1}^{n}V^{i}\right)=\frac{J}{I},$
@@
$W^{i}=\frac{\partial^{3}\ell(\nu)}{\partial\nu^{3}}=\frac{n}{8}\left\{\psi^{\prime\prime}\left(\frac{\nu+1}{2}\right)-\psi^{\prime\prime}\left(\frac{\nu}{2}\right)\right\}+\frac{1}{2}\sum_{i=1}^{n}\left\{-\frac{1}{\nu^{2}}+\frac{1}{(\nu+\hat{z}_{i\nu}^{2})^{2}}+\frac{\hat{z}_{i\nu}^{2}-1}{(\nu+\hat{z}_{i\nu}^{2})^{3}}\right\},$
@@
Appendix G Extra Results of robust regression simulations
G.1 Results when $p=3$
@@ G.2 Results when $p=80$
Appendix H Proof of Proper Posterior distribution for using $I_{\nu\nu}^{1/2}(\nu)$
@@
The results follows from Corollary 1.1 and the theorem 1 at page 156 of Fernández and Steel (1999). ∎

### v3

@@ 4.1 Simulation Settings
To investigate the performance of $t$ regression in the presence of outliers, we considered a linear model with true $\beta=(1,5,10)$ and $n=100$. Additionally, we also considered data with outliers under a high-dimensional setting, generated using a linear model with $p=80$, true $\beta=\boldsymbol{0}_{p}$, and $n=300$. The covariates and errors were generated from the standard normal distribution. To introduce outliers, the error distribution was contaminated with (i) $N(0,9)$, (ii) $\chi^{2}(4)-4$, (iii) $t(2)$, or (iv) two-point contamination in which any response value may become 5-5 or 5 with a probability $\lambda/2$ each. The contamination rates considered are $\lambda=$ 10%, 20%, and 30%.
@@ 4.2 Results when $p=3$
When the data is contaminated with $N(0,9)$ errors (see Figure 6) , although the $\nu$-estimation approaches do not necessarily give the lowest overall RMSE, they are able to give relatively low RMSE compared to OLS and Huber regression. As the contamination rate increases, the $\nu$-estimation approaches outperform most of the fixed $\nu$ approach estimations as they are successful in estimating $\nu$ close to the truth. A similar pattern is shown in the data with two-point contamination.
@@
In cases where data is contaminated with $\chi^{2}(4)-4$ (see Figure 7) or $t(2)$ errors, the $\nu$-estimation approaches outperform OLS and Huber regression. Regardless of the contamination rates, a moderate range of fixed $\nu$ values gives slightly lower overall RMSE than all $\nu$-estimation approaches.
@@ 4.3 Results when $p=80$
In the high-dimension setting where the data is contaminated with $N(0,9)$ (see Figure 8), $\chi^{2}(4)-4$, or $t(2)$ errors, although $\hat{\nu}_{adj}$ approach does not necessarily give the lowest overall RMSE, it is able to give relatively lower RMSE compared to OLS and Huber regression. This may be explained by the fact that $\hat{\nu}_{adj}$ are closer to the truth compared to other $\nu$-estimation methods (and fixing large $\nu$) in this high-dimension setting.
@@
For data contaminated with two-point errors (see Figure 9), the $\hat{\nu}_{adj}$ approach outperforms Huber and OLS when the contamination rate is low (e.g. 10%). As the contamination rate increases, Huber regression performs better as $\hat{\nu}_{adj}$ is overestimated, and hence the corresponding $\beta$ estimation performance is similar to fixing $\nu=30$. Nonetheless, the $\nu$-estimation approaches outperform OLS when the contamination rate is 20%. As the contamination rate increases to 30%, all $\nu$-estimation approaches perform as badly as most fixed-$\nu$ approaches and OLS. This may be caused by high contamination with outliers that do not follow the heavy-tailed pattern, which the $t$ distribution fails to capture.
These results collectively suggest that the $\nu$-estimation approaches, and $\hat{\nu}_{adj}$ in some high-dimension settings, offer a robust performance across various contamination types and levels, often outperforming classical regression methods.
@@ 5 Conclusions and Discussion
In this paper, we studied robust regression with Student’s $t$ error and addressed the commonly asked question on how to choose the degrees of freedom parameter in Student’s $t$ distribution in practice. We compared four frequentist and Bayesian methods in estimating the degrees of freedom $\nu$, and studied how these choices propagate to the estimation of the regression coefficient $\beta$ in uncontaminated and contaminated data. In terms of $\beta$ estimation performance, we suggest estimating $\nu$ with adjusted profile likelihood when the $p/n$ ratio is not too large, as it is the most robust in both heavy-tailed and light-tailed distributions, as well as in contaminated data and large $p$ (given moderate $n$) settings. However, to ensure stable $\nu$ and $\beta$ estimations, our proposed $\nu$-estimation method requires $p$ to be not too large relative to $n$. This is because high dimensionality may lead to larger residuals magnitude and tend to overestimate $\nu$. Therefore, when the $p/n$ ratio is very large, optimizing the likelihood with $\nu$ fixed at a large value may be more efficient. We leave a more precise characterization of this threshold for future work.
Appendix A Derivation of observed Fisher information for $\beta$ and $\sigma$
@@
Appendix B Proof of Invariance Property
@@ B.1 Profile likelihood
$l^{P}(\nu;y^{\prime},x)$ $=C(\nu)-\frac{\nu+1}{2}\sum_{i=1}^{n}\log\{\nu+z_{i}^{2}\}-n\log(a\hat{\sigma}_{\nu})=l^{P}(\nu;y,x)-n\log(a),$
@@
B.2 Adjusted Profile likelihood
@@
B.3 Full Independence Jeffreys
@@ B.4 Marginal Jeffrey’s
Appendix C More details on $\nu$-optimization
C.1 Out of bound values
@@
C.2 Choice of sample sizes and starting values
Fonseca et al. (2008) noted that, as a function of $\nu$, the profile log-likelihood function will increase without bound when $\sum_{i=1}^{n}(\hat{z_{i}}^{2}-1)^{2}<2n$ where $\hat{z_{i}}=(y_{i}-x_{i}^{T}\hat{\beta)}/\hat{\sigma}$, $\hat{\beta},\hat{\sigma}$ are the OLS estimators. Hence, to ensure that the optimization algorithm would converge, we first searched for a large enough $n$ such that Fonseca et al. (2008)’s condition is violated. Next, we searched for a sample size $n$ such that all 5 objective functions do not have flatness issues by manually plotting out the functions for 20 simulated datasets. When we ran simulations using smaller $n$, for instance, $\nu=2$; $p=40$ using $n=100$, around 3.33% of the simulations had flatness issues, and over 20% $\hat{\nu}_{adj}$ from the remaining successful converged simulations were greater than $10$, which are bad estimates. When $p$ was further increased to $60$, the flatness issues appeared in 55% of the simulations, and over 18% $\hat{\nu}_{adj}$ from the remaining successfully converged simulations were greater than $10$.
@@ C.2 Choice of sample sizes and starting values
When there are no flatness issues (i.e., the simulation settings we consider), setting the initial value at true $\omega$ that was used to generate the data optimizes the optimization time since we expect the returned $\hat{\omega}$ to be close to the truth. In our simulation settings, using a non-truth value as an initial guess would still yield stable estimates and RMSE results under the adjusted likelihood approach. The other 4 approaches tend to underestimate $\nu$, such phenomenon becomes more obvious as we set $\omega_{\text{init}}$ to be lower than the truth. For the purpose of comparing the performances between the 5 approaches, the overall RMSE conclusions were consistent with results obtained from setting the initial guess at truth, in which we concluded that the adjusted likelihood approach resulted in the lowest RMSE. For example, when we ran simulations on $n=300$; $\nu=2$; $p=60$ using the $1/5$ as initial guess of $\omega$, RMSE of $\hat{\nu}_{adj}$ is $0.5824363$, which is very close to that from setting initial guess at $1/2$ (i.e., truth). The other approaches yield RMSE at around 1.361.391.36-1.39, which are slightly inflated from setting the initial guess at the truth, which gave 1.251.291.25-1.29 RMSE. If the initial guess of $\omega$ is set at $1/10$, RMSE of $\hat{\nu}_{adj}$ is $0.5860144$, and the other approaches yield RMSE at around 1.441.491.44-1.49. The boxplots below display the spread of $\hat{\nu}$ in each setting.
@@
Appendix D Additional simulation results for $\nu$ estimation
@@
Appendix E Order of Bias and Variance for $\hat{\nu}$
@@
For the profile log-likelihood estimation of $\hat{\nu}$, both the variance and the bias have order $\mathcal{O}(n^{-1})$ asymptotically.
@@
From Reid (2013), asymptotically,
$j_{p}^{1/2}(\hat{\nu})\,(\hat{\nu}-\nu)\;\xrightarrow{\mathcal{L}}\;\mathcal{N}(0,1),$ (9)
@@
Define $f(\nu,\hat{\lambda}_{\nu})=\partial\ell(\nu,\hat{\lambda}_{\nu};y,x)/\partial\hat{\lambda}_{\nu}=0:=\ell_{\lambda}$. If $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\hat{\lambda}_{\nu}$ is invertible, taking the derivative and set $\partial f(\nu,\hat{\lambda}_{\nu})/\partial\nu=0$
$\frac{\partial\hat{\lambda}_{\nu}}{\partial\nu}=-\left(\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\hat{\lambda}_{\nu}}\right)^{-1}\frac{\partial f(\nu,\hat{\lambda}_{\nu})}{\partial\nu}.$
@@
$j_{p}(\nu)=-\ell_{\nu\nu}+\ell_{\nu\lambda}(\ell_{\lambda\lambda})^{-1}\ell_{\lambda\nu}.$
@@
The term $\ell_{\lambda\lambda}$ is the observed Fisher information sub-matrix in the adjusted profile likelihood, $j_{\lambda\lambda}$, which has order of $\mathcal{O}(n)$. Therefore, $j_{p}(\nu)=\mathcal{O}(n)$, so we have the $Var(\hat{\nu})=\mathcal{O}(n^{-1})$.
Next, we analyze the bias part; we assume the maximum likelihood estimates of $\beta,\sigma$, and $\nu$ exist. The maximum profile likelihood estimates coincide with the maximum likelihood estimates, so it suffices to show the order of bias for the maximum likelihood estimate of $\nu$. We follow the proof by Cox and Snell (1968). The maximum likelihood equation $\ell^{\prime}(\hat{\nu},\hat{\beta},\hat{\sigma};y,x)=0$ to the first order for $\nu$ is
@@
$U^{i}=\frac{\partial\ell(\nu,\hat{\beta},\hat{\sigma};y_{i},x_{i})}{\partial\nu},V^{i}=\frac{\partial^{2}\ell(\nu,\hat{\beta},\hat{\sigma};y_{i},x_{i})}{\partial\nu^{2}},$
@@
To obtain a more refined answer, we replace it with the second-order equation
@@
$cov(\hat{\nu}-\nu,\ell^{\prime\prime}(\nu))=cov\left(\frac{\sum_{i=1}^{n}U^{i}}{I},\sum_{i=1}^{n}V^{i}\right)=\frac{1}{I}cov\left(\sum_{i=1}^{n}U^{i},\sum_{i=1}^{n}V^{i}\right)=\frac{J}{I},$
@@
$W^{i}=\frac{\partial^{3}\ell(\nu,\hat{\beta},\hat{\sigma};y_{i},x_{i})}{\partial\nu^{3}}=\frac{1}{8}\left\{\psi^{\prime\prime}\left(\frac{\nu+1}{2}\right)-\psi^{\prime\prime}\left(\frac{\nu}{2}\right)\right\}+\frac{1}{2}\left\{-\frac{1}{\nu^{2}}+\frac{1}{(\nu+\hat{z}_{i\nu}^{2})^{2}}+\frac{\hat{z}_{i\nu}^{2}-1}{(\nu+\hat{z}_{i\nu}^{2})^{3}}\right\},$
@@
Appendix F Additional simulation results for robust regression
F.1 Results when $p=3$
@@ F.2 Results when $p=80$
Appendix G Proof of Proper Posterior distribution for using $I_{\nu\nu}^{1/2}(\nu)$
@@
The results follow from Corollary 1.1 and Theorem 1 on page 156 of Fernández and Steel (1999). ∎